OpenStudy (anonymous):

what is the simplification of the expression (−2x + 4y)(3x − 7y).

OpenStudy (ghazi):

use distributive law

OpenStudy (ghazi):

(-2x*3x + -2x*-7y + 4y*3x + 4y*-7y)

OpenStudy (cwrw238):

can you simplify this any further deja?

OpenStudy (cwrw238):

-2x*3x = -2*3*x*x = -6x^2 - do u understand that bit?